Two boys trading baseball cards agree that 6 Will Clark cards and 3 Jose Canseco cards are worth the same to them as 2 Nolan Ryan cards. They also agree that 9 Canseco cards are worth the same as 12 Clark cards plus 2 Ryan cards. How many Canseco cards have the same value to them as 2 Ryan cards.
Can someone help me set this one up and solve it? Thank you.

Answer by uma(370)

Let W be Will Clark cards , J be Jose Canseco cards and N be Nolan Ryan cards
Given that 6W + 3J = 2N ----------(1)
and 12W + 2N = 9J ----------------(2)
multiplying (1) by 2 gives 12W + 6J = 4N -----------------(3)
12W - 9J = - 2N ---------------(2) rearranged
Subtracting the above two equations we have, 6J + 9J = 4N + 2N
==> 15J = 6N
==> 5J = 2N [Dividing both the sides by 3]
Therefore 2 Ryan cards = 5 Canseco cards